Consider the following data.

2,2,2,−7,−7,−6,−72,2,2,−7,−7,−6,−7

Copy Data

Step 1 of 3:

Determine the mean of the given data.
Answer #1


Solution

The sample mean is

Mean = (x / n) )

=2+2+2+-7+-7+-6+-72+2+2+-7+-7+-6+-7 /13

=-109 /13

=-8.3846

Mean = 8.38